the Time between two arrivals in a queuing model is normally distributed with mean 2 minutes and standard deviation 0.25 minutes if random samples of size 36 are drawn what is the probability that the sample mean will be less then 2.25 minutes ?


1
Expert's answer
2021-06-22T14:21:22-0400

"\mu=2 \\
\sigma=0.25 \\
n=36 \\
P(\bar{x}<2.25) = P(Z< \frac{2.25-2}{0.25/ \sqrt{36}}) \\
= P(Z<6) \\
= 0.9999"